the value of lim x tends to 0 ex^2 -cos x / x2 is.....................................
 
 

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
We can apply L hospitals rule in this problem too
             Lim ( x à0) e^x^2- Cos x /x^2
            = lim(xà0)[ 2xe^x  + Sin x]/ 2x
            = Lim (xà0) [ 4x^2 e^x^2 + 2 e^x^2 + Cos x ]/2
            = (0+2+1)/2
            = 3/2.
Final limit value is 1.5.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)